OpenStudy (anonymous):

What is the slope of the line perpendicular to segment EG if the coordinates of E and G are (-3, 2) and (2, 0)? 5/2 -2/5 -5/2 2/5

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

First find the slope of EG. We can use the slope formula \(m = \dfrac{y_2-y_1}{x_2-x_1}\) Plug in the two points: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{0-2}{2+3}\) Add & Subtract: \(m = -\dfrac{2}{5}\) So the slope of EG is -2/5. Perpendicular lines have opposite slope so what's the opposite of -2/5? @puppylover234

OpenStudy (anonymous):

5/2?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yep! you got it!

OpenStudy (anonymous):

thanks a ton! I really appreciate the help! you rock!

OpenStudy (igreen):

No problem! xD You can write fan me and write a testimonial if you want.

OpenStudy (igreen):

@puppylover234